Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form, ax^2 + bx + c = 0
Solution 6, only solution
The equation is ??? = 0

Answer by mananth(16946)

since 6 is the solution
x=6
x-6=0
(x-6)^2=0
x^2-12x+36=0
...
m.ananth@hotmail.ca